OpenStudy (anonymous):

solve the equation -2v - 7 = -23

OpenStudy (anonymous):

Add 7 to both sides, then divide both sides by -2.

OpenStudy (anonymous):

so how would i do that

OpenStudy (anonymous):

-2v - 7 = -23 Add 7 to both sides: -2v - 7 +7 = -23 + 7 Can you simplify it?

OpenStudy (anonymous):

v = 8 right

OpenStudy (anonymous):

Yes! :)